### v1

@@ Acknowledgment
The authors are grateful to Ivan Seskar and ORBIT staff for setting up the radios on ORBIT. The authors thank Loukas Lazos, Ming Li, Jingcheng Li, Ziqi Xu, Samuel H. Knarr, and Timothy C. Burt for valuable advice, and sharing the equipment. Finally, the authors acknowledge helpful discussions with Mark J. Meisner, Jaim Bucay, Dennis L. Goeckel, Donald F. Towsley, Matthieu R. Bloch, Matthew Arcarese, and Robert J. McGurrin.
@@ Appendix A Covertness Criterion Analysis
$\psi(\vec{w};\vec{\mu},\mathbf{\Gamma},\mathbf{C})={\frac{\exp\!\left(-{\frac{1}{2}}{\begin{bmatrix}\vec{w}-\vec{\mu}\\ {\overline{\vec{w}}}-{\overline{\vec{\mu}}}\end{bmatrix}}^{\dagger}{\begin{bmatrix}\mathbf{\Gamma}&\mathbf{C}\\ {\overline{\mathbf{C}}}&{\overline{\mathbf{\Gamma}}}\end{bmatrix}}^{\!\!-1}\!{\begin{bmatrix}\vec{w}-\vec{\mu}\\ {\overline{\vec{w}}}-{\overline{\vec{\mu}}}\end{bmatrix}}\right)}{\pi^{n}{\sqrt{\det(\mathbf{\Gamma})\det(\overline{\mathbf{\Gamma}}-\mathbf{C}^{\dagger}\mathbf{\Gamma}^{-1}\mathbf{C})}}}}.$ (14)
Since our noise is i.i.d. and induces circularly symmetric complex Gaussian distributions, $\mathbf{C}=\mathbf{0}$ and we drop it for brevity. Under $H_{0}$, Alice does not transmit and Willie observes AWGN. Thus, segments are i.i.d. with the density function:
@@
$\phantom{=}+\frac{\alpha_{n}}{4}\psi\left(\vec{w}_{i};\vec{0},2\sigma^{2}_{w}\mathbf{I}_{n_{s}}\right)e^{-a_{w}^{2}\frac{\|\vec{c}_{p}\|^{2}+\|\vec{c}_{q}\|^{2}}{2\sigma_{w}^{2}}}$ (18)
@@
$H^{2}$ $(P_{0}^{n},P_{1}^{n})$
@@
$\delta\leq\sqrt{\frac{1}{2}\left(1-\left(1-\alpha_{n}^{2}\frac{1+r_{p/q}^{4}}{32\sigma_{w}^{4}}a_{w}^{4}c_{q}^{4}\right)^{n_{p}}\right)}$ (26)
@@
Then, for $n_{p}$ large enough, Taylor’s theorem with remainder and reverting the substitution implies
$\alpha_{n}\leq\frac{4\sigma_{w}^{2}}{a_{w}^{2}c_{q}^{2}}\sqrt{\frac{2\log\left(\frac{1}{1-2\delta^{2}}\right)}{n_{p}(1+r_{p/q}^{4})}}.$ (29)
@@ Appendix B Phase Estimation
since AWGN is circularly-symmetric and has zero mean. The estimate of phase $\theta_{b}$ is then $\hat{\theta}_{b}=\tan^{-1}\frac{p_{I}}{p_{Q}}$. As $n_{s}^{(p)}$ increases, $p_{I}$ and $p_{Q}$ converge to their expected values, yielding a more accurate estimate of phase. Here, $n_{s}^{(p)}=6$ suffices.
Appendix C Estimation of Willie’s SNR
We need to estimate Willie’s SNR given the knowledge of $\vec{c}_{p}$, $\vec{c}_{q}$, as well as transmitted symbols in $\vec{x}_{n_{t}}$ and their locations $\vec{t}$. While Willie has no access to $\vec{x}_{n_{t}}$ and $\vec{t}$, we use them to characterize the SNR of his system. A pulse-bearing slot received by Willie is:
$\vec{w}_{p}(x)$ $=a_{w}e^{j\theta_{w}}\vec{u}(x)+\vec{z}^{(w)},$ (34)
for $x\in\mathcal{X}$ and unknown phase $\theta_{w}$. Applying the pulse shape $\vec{u}(x)$ corresponding to transmitted $x$ as a filter yields:
$\langle\vec{u}(x),\vec{w}_{p}(x)\rangle$ $=a_{w}e^{j\theta_{w}}c^{2}+\langle\vec{u}(x),\vec{z}^{(w)}\rangle,$ (35)
since $\|\vec{u}(x)\|=c$ per Section 2. The expected value of the above is:
$\mathbb{E}\left[\langle\vec{u}(x),\vec{w}_{p}(x)\rangle\right]$ $=a_{w}e^{j\theta_{w}}c^{2},$ (36)
as AWGN has zero mean, and, hence, $\mathbb{E}\left[\langle\vec{u}(x),\vec{z}^{(w)}\rangle\right]=0$. Thus, averaging over many pulse-bearing slot observations, and dividing by a known constant $c^{2}$, yields an estimate of $a_{w}e^{j\theta_{w}}$. Its squared magnitude is the estimate $\hat{a}_{w}^{2}$ of $a_{w}^{2}$.
@@ Appendix C Estimation of Willie’s SNR
An empty pulse slot received by Willie contains only noise:
@@
Thus, averaging over many instances of empty pulse slots, and dividing by a known constant $2n_{s}$, yields an estimate $\hat{\sigma}_{w}^{2}$ of $\sigma_{w}^{2}$. Finally, we estimate the SNR by evaluating $\frac{\hat{a}_{w}^{2}c_{q}^{2}}{\hat{\sigma}_{w}^{2}}$, for known $c_{q}=\|\vec{c}_{q}\|$.

### v2

@@ Acknowledgment
The authors are grateful to Ivan Seskar and the COSMOS staff for setting up the radios on COSMOS and answering our many questions. The authors thank Loukas Lazos, Ming Li, Jingcheng Li, Ziqi Xu, Samuel H. Knarr, and Timothy C. Burt for valuable advice and for sharing the equipment. The authors also acknowledge helpful discussions with Mark J. Meisner, Jaim Bucay, Dennis L. Goeckel, Donald F. Towsley, Matthieu R. Bloch, Matthew Arcarese, and Robert J. McGurrin. Finally, the authors thank the anonymous referees for their suggestions, which substantially improved the manuscript.
@@ Appendix A Covertness Criterion Analysis
$\psi(\vec{w};\vec{\mu},\mathbf{\Gamma},\mathbf{C})={\frac{\exp\!\left(-{\frac{1}{2}}{\begin{bmatrix}\vec{w}-\vec{\mu}\\ {\overline{\vec{w}}}-{\overline{\vec{\mu}}}\end{bmatrix}}^{\dagger}{\begin{bmatrix}\mathbf{\Gamma}&\mathbf{C}\\ {\overline{\mathbf{C}}}&{\overline{\mathbf{\Gamma}}}\end{bmatrix}}^{\!\!-1}\!{\begin{bmatrix}\vec{w}-\vec{\mu}\\ {\overline{\vec{w}}}-{\overline{\vec{\mu}}}\end{bmatrix}}\right)}{\pi^{n}{\sqrt{\det(\mathbf{\Gamma})\det(\overline{\mathbf{\Gamma}}-\mathbf{C}^{\dagger}\mathbf{\Gamma}^{-1}\mathbf{C})}}}},$ (17)
where $\overline{\mathbf{A}}$ denotes the complex conjugate of matrix $\mathbf{A}$. Since our noise is i.i.d. and induces circularly symmetric complex Gaussian distributions, $\mathbf{C}=\mathbf{0}$ and we drop it for brevity. Under $H_{0}$, Alice does not transmit and Willie observes AWGN. Thus, segments are i.i.d. with the density function:
@@
$\phantom{=}+\frac{\alpha_{n}}{4}\psi\left(\vec{w}_{i};\vec{0},2\sigma^{2}_{w}\mathbf{I}_{n_{s}}\right)e^{-a_{w}^{2}\frac{\|\vec{c}_{p}\|^{2}+\|\vec{c}_{q}\|^{2}}{2\sigma_{w}^{2}}}$
@@
$\int_{\mathbb{C}^{n_{s}}}$ $\mathop{}\!\mathrm{d}^{2}\vec{w}_{1}\sqrt{p^{(0)}\left(\vec{w}_{1}\right)p^{(1)}\left(\vec{w}_{1}\right)}$
@@
$\delta$ $\geq\sqrt{\frac{1}{2}\left(1-\left(1-\alpha_{n}^{2}\frac{1+r_{p/q}^{4}}{32\sigma_{w}^{4}}a_{w}^{4}c_{q}^{4}\right)^{n_{p}}\right)}$ (26)
@@
Then, for $n_{p}$ large enough, Taylor’s theorem with remainder and reverting the substitution yields the expression in (13).
@@ Appendix B Phase Estimation
since AWGN is circularly-symmetric and has zero mean. The estimate of phase $\theta_{b}$ is then $\hat{\theta}_{b}=\tan^{-1}\frac{p_{I}}{p_{Q}}$. As $n_{s}^{(p)}$ increases, $p_{I}$ and $p_{Q}$ converge to their expected values, yielding a more accurate estimate of phase. Here, $n_{s}^{(p)}=26$ suffices.
Appendix C Derivation of Optimal Detector Test Statistic
As in Appendix A, we substitute $\vec{c}_{q}\to c_{q}\vec{b}_{q}$ and $\vec{c}_{p}\to c_{q}r_{p/q}\vec{b}_{p}$ in the expression for $L_{i}$ in (14) and Taylor expand it about $c_{q}=0$:
$L_{i}$ $=\frac{\alpha_{n}}{4\sigma_{w}^{4}}\left(r_{p/q}^{2}\left(|\langle\vec{b}_{p},\vec{p}_{i}\rangle|^{2}-2\sigma_{w}^{2}||\vec{b}_{p}||^{2}\right)\right$.
[…]
Thus, for $c_{q}$ sufficiently small, the $i^{\text{th}}$ pulse-slot component $S_{i}^{\text{(opt)}}$ of Willie’s sufficient statistic is in (15).
Appendix D Willie’s Detectors’ Errors
D-A Analysis
Here, we derive bounds for the probability of error attainable by the optimal detector and radiometer, described in Section II-E. We use the Berry-Esseen theorem and assume non-informative priors. Both test statistics $S^{\text{(opt)}}(\vec{w})$ and $S^{\text{(rad)}}(\vec{w})$ have the form $S=\sum_{i=1}^{n_{p}}S_{i}$, where $(S_{i})$ is a sequence of realizations of $n_{p}$ i.i.d. random variables. Let us standardize this sum using $H_{0}$ parameters:
$S_{n_{p}}=\frac{S-n_{p}\mu_{0}}{\sigma_{0}\sqrt{n_{p}}},$ (34)
where $\mu_{0}$ and $\sigma_{0}^{2}$ are the mean and variance of $S_{i}$ under $H_{0}$. Similarly, let $\mu_{1}$ and $\sigma_{1}^{2}$ be the mean and variance of $S_{i}$ under $H_{1}$, noting that they may depend on $n_{p}$. We provide the necessary moments of $S_{i}$ under $H_{0}$ and $H_{1}$ for both optimal detector and radiometer in Appendix E.
Let $F_{n_{p},h}(x)$ be the cumulative distribution function (c.d.f.) of the random variable statically describing $S_{n_{p}}$ under $H_{h}$, $h\in\{0,1\}$. By the Berry-Esseen theorem [22, Ch. XVI.5],
$\sup_{x}|F_{n_{p},k}(x)-\Phi(x)|\leq B_{n_{p},k}\triangleq\frac{C\rho_{k}}{\sigma_{k}^{3}\sqrt{n_{p}}},$ (35)
where $\rho_{k}$ and $\sigma_{k}$ are the absolute third central moment and the standard deviation of $S_{i}$ under $H_{k}$, $\Phi(x)$ is the c.d.f. of standard Gaussian distribution $\mathcal{N}(0,1)$, and $C<0.4748$ is the Berry-Esseen constant. Willie decides between $H_{0}$ and $H_{1}$ by comparing $S_{n_{p}}$ to a threshold $\tau$, per Section II-E. The probability of error given $\tau$ is $p_{e}^{(w)}(\tau)=\frac{1}{2}\left(P(S_{n_{p}}>\tau|H_{0})+P(S_{n_{p}}\leq\tau|H_{1})\right)$. Applying (35), we have $P(S_{n_{p}}>\tau|H_{0})=Q(\tau)+\epsilon_{n_{p},0}$, where the approximation error $\epsilon_{n_{p},0}\in[-B_{n_{p},0},B_{n_{p},0}]$ and $Q(x)\triangleq\frac{1}{\sqrt{2\pi}}\int_{x}^{\infty}e^{-t^{2}/2}\mathop{}\!\mathrm{d}t$ is the Gaussian error function. Under $H_{1}$, we transform the event $S_{n}\leq\tau$ to the standardized frame of $H_{1}$. Defining the effective argument $z(\tau)\triangleq\frac{\tau\sigma_{0}-\sqrt{n_{p}}\Delta\mu}{\sigma_{1}}$, where $\Delta\mu=\mu_{1}-\mu_{0}$, we have $P(S_{n_{p}}\leq\tau|H_{1})=Q\left(-z(\tau)\right)+\epsilon_{n_{p},1}$, where $\epsilon_{n_{p},1}\in[-B_{n_{p},1},B_{n_{p},1}]$. Thus, $p_{e}^{(w)}(\tau)=\frac{1}{2}\left(Q(\tau)+Q\left(-z(\tau)\right)\right)+\frac{\epsilon_{n_{p},0}+\epsilon_{n_{p},1}}{2}$. The sum of $Q$-functions is minimized by symmetric arguments: $\tau=-z(\tau)=\frac{\sqrt{n_{p}}\Delta\mu}{\sigma_{0}+\sigma_{1}}$. Thus, the minimum error probability achievable by both the optimum detector and the radiometer is:
$p_{e,\text{det}}^{(w)}$ $=Q\left(\frac{\sqrt{n_{p}}\Delta\mu}{\sigma_{0}+\sigma_{1}}\right)+\frac{\epsilon_{n_{p},0}+\epsilon_{n_{p},1}}{2},$ (36)
with the distinction arising from the expressions for the moments being different for the detectors.
D-B Discussion
When Alice follows the SRL, for both detectors, per Appendix E, $\Delta\mu\propto\frac{1}{\sqrt{n_{p}}}$ and $\sigma_{1}\to\sigma_{0}$ as $n_{p}\to\infty$. Thus, $p_{e,\text{det}}^{(w)}$ converges to a constant, meaning neither detector is effective. On the other hand, when Alice does not follow the SRL, $p_{e,\text{det}}^{(w)}\to 0$ as $n_{p}\to\infty$, since $\sigma_{0}+\sigma_{1}$ are upper-bounded by a constant, while $\sqrt{n_{p}}\Delta\mu$ is unbounded and $\epsilon_{n_{p},0},\epsilon_{n_{p},1}\to 0$. This is borne out in our experiments.
Furthermore, since $p_{e,\text{det}}^{(w)}$ converges to $p_{e}^{(w)}$ in (10) for optimal detector as $n_{p}\to\infty$, Taylor series expansion of (36) around $\alpha_{n}=0$ yields (16), which translates into $\approx 25\%$ improvement in $\alpha_{n}$ for $\delta=0.07$ used here.
Appendix E Moments of Willie’s Detectors’ Sufficient Statistics
To analyze the error probability, we require the mean, variance, and upper bounds on the third absolute central moment of $S_{i}^{\text{(opt)}}$ and $S_{i}^{\text{(rad)}}$ under both hypotheses.
E-A Optimal Detector
Under $H_{0}$, observations follow complex Gaussian distributions $\vec{w}_{i}=(\vec{p}_{i},\vec{q}_{i})\sim\mathcal{CN}(\vec{0},2\sigma_{w}^{2}\mathbf{I}_{n_{s}})$. Consequently, $S_{i}^{\text{(opt)}}$ is a sum of two independent exponential random variables. The mean $\mu_{0}$ and variance $\sigma_{0}^{2}$ are
$\mu_{0}$ $=2\sigma_{w}^{2}\left(\|\vec{c}_{p}\|^{2}+\|\vec{c}_{q}\|^{2}\right),$ (37)
$\sigma_{0}^{2}$ $=4\sigma_{w}^{4}\left(\|\vec{c}_{p}\|^{4}+\|\vec{c}_{q}\|^{4}\right).$ (38)
The Berry-Esseen bound relies on the third absolute central moment $\rho_{0}=\mathbb{E}\left[|S_{i}-\mu_{0}|^{3}\right]$. We write $S_{i}-\mu_{0}=(X-\mu_{x})+(Y-\mu_{y})$, where $X$ and $Y$ are independent exponential variables corresponding to the pilot and QPSK projections. Using the Minkowski inequality for the $L_{3}$ norm
$\rho_{0}$ $=\|(X-\mu_{x})+(Y-\mu_{y})\|_{3}^{3}$ (39)
$\leq\left(\|X-\mu_{x}\|_{3}+\|Y-\mu_{y}\|_{3}\right)^{3}.$ (40)
For an exponential variable $X$, $\|X-\mu_{x}\|_{3}=\mu_{x}(12/e-2)^{1/3}$. Substituting this yields an upper bound for $\rho_{0}$
$\rho_{0}\leq\left(\frac{12}{e}-2\right)(\mu_{x}+\mu_{y})^{3}=\left(\frac{12}{e}-2\right)\mu_{0}^{3}\approx 2.415\mu_{0}^{3}.$ (41)
Under $H_{1}$, Willie receives a statistical mixture: with probability $1-\alpha_{n}$, the slot contains only noise ($t=0$), and with probability $\alpha_{n}$, it contains signal ($t=1$). The conditional mean given $t=1$ is shifted by the signal energy $\Delta P=a_{w}^{2}\left(\|\vec{c}_{p}\|^{4}+\|\vec{c}_{q}\|^{4}\right)$
$\mu_{P}=\mathbb{E}[S_{i}|t=1]$ $=\mu_{0}+\Delta P.$ (42)
The unconditional mean under $H_{1}$ is $\mu_{1}=\mu_{0}+\alpha_{n}(\mu_{P}-\mu_{0})=\mu_{0}+\alpha_{n}\Delta P$. Note that the difference in means is given by $\Delta\mu=\alpha_{n}\Delta P$.
The conditional variance given $t=1$ involves non-central terms
$\sigma_{P}^{2}=\mathrm{Var}[S_{i}|t=1]$ $=\sigma_{0}^{2}+4\sigma_{w}^{2}a_{w}^{2}\left(\|\vec{c}_{p}\|^{6}+\|\vec{c}_{q}\|^{6}\right).$ (43)
Using the law of total variance, the unconditional variance $\sigma_{1}^{2}$ is
$\sigma_{1}^{2}$ $=\sigma_{0}^{2}+\alpha_{n}(\sigma_{P}^{2}-\sigma_{0}^{2})+\alpha_{n}(1-\alpha_{n})(\Delta P)^{2}.$ (44)
Analytically evaluating the third absolute central moment $\rho_{1}$ for the mixture distribution is intractable due to the absolute value function. Instead, we bound $\rho_{1}$ using the fourth central moment $\mu_{4,1}$ via Lyapunov’s inequality $\rho_{1}\leq\left(\mu_{4,1}\right)^{3/4}$. The fourth central moment is derived algebraically from the raw moments. Let $m_{k}^{(0)}=\mathbb{E}\left[S_{i}^{k}\middle|t=0\right]$ and $m_{k}^{(1)}=\mathbb{E}\left[S_{i}^{k}\middle|t=1\right]$. The mixture raw moments are
$\mathbb{E}[S_{i}^{k}]=(1-\alpha_{n})m_{k}^{(0)}+\alpha_{n}m_{k}^{(1)}.$ (45)
From these, $\mu_{4,1}$ is obtained via the expansion
$\mu_{4,1}=\mathbb{E}[S_{i}^{4}]-4\mu_{1}\mathbb{E}[S_{i}^{3}]+6\mu_{1}^{2}\mathbb{E}[S_{i}^{2}]-3\mu_{1}^{4}$ (46)
and raw moments [28, Eq. following (29.31)]
$m_{k}^{(1)}$ $=\sum_{j=0}^{k}\binom{k}{j}\left[j!\beta_{p}^{j}L_{j}\left(-\frac{\lambda_{p}}{\beta_{p}}\right)\right]$ (47)
$\phantom{======}\times\left[(k-j)!\beta_{q}^{k-j}L_{k-j}\left(-\frac{\lambda_{q}}{\beta_{q}}\right)\right],$ (48)
where $L_{k}(u)$ are Laguerre polynomials and $\beta_{p}=2\sigma_{w}^{2}\|\vec{c}_{p}\|^{2}$ and $\lambda_{p}=a_{w}^{2}\|\vec{c}_{p}\|^{4}$ denote the noise power and signal energy for the pilot (similarly $\beta_{q}=2\sigma_{w}^{2}\|\vec{c}_{q}\|^{2},\lambda_{q}=a_{w}^{2}\|\vec{c}_{q}\|^{4}$ for QPSK).
E-B Power Detector
Under $H_{0}$, $\vec{w}_{i}\sim\mathcal{CN}(\vec{0},2\sigma^{2}_{w}\mathbf{I}_{n_{s}})$. Consequently, $S_{i}^{\text{(rad)}}=\|\vec{w}_{i}\|^{2}$ is a scaled central $\chi^{2}$ random variable with $2n_{s}$ degrees of freedom and scaling factor $\sigma_{w}^{2}$. The mean and variance are
$\mu_{0}$ $=2n_{s}\sigma_{w}^{2},$ (49)
$\sigma_{0}^{2}$ $=4n_{s}\sigma_{w}^{4}.$ (50)
We again bound the third absolute central moment $\rho_{0}$ via Lyapunov’s inequality $\rho_{0}\leq(\mu_{4,0})^{3/4}$. The fourth central moment is given by
$\mu_{4,0}$ $=48n_{s}\sigma_{w}^{8}(n_{s}+2).$ (51)
Under $H_{1}$, the detector statistic is drawn from a mixture distribution. Conditioned on signal presence ($t=1$), $S_{i}^{\text{(rad)}}$ follows a scaled non-central $\chi^{2}$ distribution with $2n_{s}$ degrees of freedom and non-centrality parameter $\lambda=\frac{\Delta P}{\sigma_{w}^{2}}$, with $\Delta P=a_{w}^{2}\left(\|\vec{c}_{p}\|^{2}+\|\vec{c}_{q}\|^{2}\right)$. The conditional mean is then $\mu_{P}=\mu_{0}+\Delta P$, and the unconditional mean is given by
$\mu_{1}$ $=\mu_{0}+\alpha_{n}\Delta P.$ (52)
The conditional variance is $\sigma_{P}^{2}=\sigma_{0}^{2}+4\sigma_{w}^{2}\Delta P$. Applying the law of total variance, we obtain
$\sigma_{1}^{2}$ $=\sigma_{0}^{2}+\alpha_{n}(\sigma_{P}^{2}-\sigma_{0}^{2})+\alpha_{n}(1-\alpha_{n})\Delta P^{2}$ (53)
$=\sigma_{0}^{2}+4\alpha_{n}\sigma_{w}^{2}\Delta P+\alpha_{n}(1-\alpha_{n})\Delta P^{2}.$ (54)
We find the fourth central moment $\mu_{4,1}$ by expanding the moments of the mixture distribution. Let $\Delta\mu=\mu_{1}-\mu_{0}=\alpha_{n}\Delta P$ and $\Delta\mu^{\prime}=\mu_{P}-\mu_{1}$. Now, we have
$\mu_{4,1}$ $=(1-\alpha_{n})\mathbb{E}\left[\left(S_{i}^{\text{(rad)}}-\mu_{0}-\Delta\mu\right)^{4}\middle|t=0\right]$
$\phantom{=}+\alpha_{n}\mathbb{E}\left[\left(S_{i}^{\text{(rad)}}-\mu_{P}+\Delta\mu^{\prime}\right)^{4}\middle|t=1\right].$ (55)
Expanding the centered moments, the linear terms disappear, leaving
$\mu_{4,1}$ $=(1-\alpha_{n})\left[\mu_{4,0}-4\Delta\mu\mu_{3,0}+6(\Delta\mu)^{2}\sigma_{0}^{2}+(\Delta\mu)^{4}\right]$
$\phantom{=}+\alpha_{n}\left[\mu_{4,P}+4\Delta\mu^{\prime}\mu_{3,P}+6(\Delta\mu^{\prime})^{2}\sigma_{P}^{2}+(\Delta\mu^{\prime})^{4}\right].$ (56)
The third central moment for $t=0$ is $\mu_{3,0}=16n_{s}\sigma_{w}^{6}$. For $t=1$, the moments are given by the cumulants of the scaled non-central $\chi^{2}$ distribution [28, Ch. 29] as
$\mu_{3,P}$ $=8\sigma_{w}^{6}\left(2n_{s}+3\frac{\Delta P}{\sigma_{w}^{2}}\right)$ (57)
$\mu_{4,P}$ $=48\sigma_{w}^{8}\left(2n_{s}+4\frac{\Delta P}{\sigma_{w}^{2}}\right)+3\sigma_{P}^{4}.$ (58)
Appendix F Estimation of Willie’s Channel Parameters
@@
We need to estimate Willie’s channel parameters from his observation and knowledge of $\vec{c}_{p}$ and $\vec{c}_{q}$. An empty pulse slot received by Willie contains only noise:
@@
Thus, averaging over many instances of empty pulse slots and dividing by a known constant $2n_{s}$ yields an estimate $\hat{\sigma}_{w}^{2}$ of $\sigma_{w}^{2}$. This can be used to estimate $\mu_{0}$ and $\sigma_{0}^{2}$ for both the optimal detector and the radiometer, as per Appendix E. Similarly, one can estimate $a_{w}^{2}$ from the pulse slots used for transmission.
Now, denote the test statistic for either detector under hypothesis $H_{h}$, $h\in\{0,1\}$ by $S^{(h)}$. After normalizing, $\frac{S^{(h)}}{n_{p}}\to\mathbb{E}\left[S^{(h)}_{i}\right$] by Law of Large Numbers. Arithmetic and estimate of $\sigma_{w}^{2}$ yield the quantities we need to calculate the threshold.